OpenStudy (anonymous):

please help: 4^x+4 = √2^5x find the value of x.

OpenStudy (joannablackwelder):

If you can make the bases equal, you can set the exponents equal to each other.

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

how do i remove the sqared root

OpenStudy (joannablackwelder):

You substitute.

OpenStudy (joannablackwelder):

I think it is easier to rewrite 4 in terms of sqrt(2) personally.

OpenStudy (anonymous):

oh i was thinking like this √2^5x >> 2^5x/2

OpenStudy (joannablackwelder):

That works too.

OpenStudy (anonymous):

is this right 4^x+4 >> √2^x+6 ?

OpenStudy (joannablackwelder):

No.

OpenStudy (joannablackwelder):

You already got 2 as a base from the sqrt(2) side. Why not just make a base of 2 on the 4 side too?

OpenStudy (joannablackwelder):

4 = 2^2

OpenStudy (anonymous):

oh i see 4 sqrt(2) would be 2√2

OpenStudy (anonymous):

so 4^x+4 >> √2^x=4 ??

OpenStudy (anonymous):

2√2^x+4

OpenStudy (joannablackwelder):

I don't understand how or why you did that...

OpenStudy (anonymous):

im lost would you show me how to convert 4 to sqrt(2) please

OpenStudy (joannablackwelder):

You said sqrt(2)^5x ->> 2^5x/2 That was a good first step.

OpenStudy (joannablackwelder):

You can also rewrite 4^x+4 as 2^2(x+4)

OpenStudy (anonymous):

ok so it would be like this 4^x+4 = 2 5x/2 then 2.2^x+4 = 2 ^5x/2 then i operate the exponents x+4=5x/2

OpenStudy (anonymous):

so how do i find x in there

OpenStudy (joannablackwelder):

The exponents are 2(x+4) and 5x/2

OpenStudy (joannablackwelder):

When you raise an exponent to another exponent, you multiply.

OpenStudy (anonymous):

oh yeah i made a mistake 4^x+4 would be 2^2x+4

OpenStudy (anonymous):

so it would be like this then 2x+8=2^5x/2 how do i find x in there?

OpenStudy (anonymous):

i mean 2x+8=5x/2

OpenStudy (joannablackwelder):

Right.

OpenStudy (anonymous):

so now 5x=2x+8*2

OpenStudy (joannablackwelder):

Right, but distributed.

OpenStudy (joannablackwelder):

Don't forget parentheses.

OpenStudy (anonymous):

like this 5x=2x+16 5x-2x=16 3x=16 x=16/3

OpenStudy (joannablackwelder):

Distribute! 5x=4x+16

OpenStudy (joannablackwelder):

The 2 is multiplied by both terms.

OpenStudy (anonymous):

oh i seeee

OpenStudy (joannablackwelder):

:)

OpenStudy (anonymous):

so x would be 16 :|

OpenStudy (joannablackwelder):

Right! :)

OpenStudy (anonymous):

omg thanks.

OpenStudy (anonymous):

i have a 16 question test :}

OpenStudy (anonymous):

online

OpenStudy (joannablackwelder):

No worries. Just remember to distribute!

OpenStudy (anonymous):

hehe.

OpenStudy (joannablackwelder):

Test?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

on thatquiz.org

OpenStudy (anonymous):

16 question 120 min to answer im so scared

OpenStudy (joannablackwelder):

We aren 't allowed to help on tests or quizzes.

OpenStudy (anonymous):

oh that too bad/

OpenStudy (joannablackwelder):

Or are you prepping for it?

OpenStudy (anonymous):

yeah i was

OpenStudy (joannablackwelder):

That's cool. Just not during the test itself. :)

OpenStudy (joannablackwelder):

Happy studies!do you have some practice problems where you can check your answers?

OpenStudy (anonymous):

no really

OpenStudy (joannablackwelder):

Those are the best to practice with.

OpenStudy (anonymous):

jsut the question for instance this one says calculate log20 + log 5 but i dont have answer anywhere

OpenStudy (joannablackwelder):

Is this alg. 2?

OpenStudy (anonymous):

you mean algebra 2?

OpenStudy (joannablackwelder):

Yeah

OpenStudy (anonymous):

yeah that's algebra

OpenStudy (anonymous):

i have a test on algebra and geometry

OpenStudy (joannablackwelder):

There is a website called Kuta Software that has practice problems with answers for algebra and geometry.

OpenStudy (anonymous):

oh thats cool.

OpenStudy (anonymous):

is that a software?

OpenStudy (joannablackwelder):

Practicing but not knowing if your answers are right or wrong doesn 't help too much in my opinion.

OpenStudy (anonymous):

yeah i agree.

OpenStudy (joannablackwelder):

You can just download individual pdfs.

OpenStudy (joannablackwelder):

Take a look. :)

OpenStudy (anonymous):

oh nice

OpenStudy (joannablackwelder):

And, by the way, you are in the physics section. This question belongs more in the math one.

OpenStudy (anonymous):

yeah i realized that my mistake :| thanks for the website i found the pdfs:)

OpenStudy (joannablackwelder):

No worries. Happy studies! :)

OpenStudy (anonymous):

JoannaBlackwelder you are awesome. I wish I knew how to give medals or kudos. I just followed this problem out of curiosity but found it very enlightening!

OpenStudy (joannablackwelder):

Thanks, @patrickdixon ! You can give me some kudos by clicking the Best response button. Glad it helped you! :)